which is greater molality or molarity?why?

Grade:12

4 Answers

Sunil Kumar FP
10 years ago
its not about which is greater.it is simply a method of expressing concentration of a solution.
molarity is the no of mole of solute in a given volume of solution.
molality is moles of solute in a given (per kg) of solvent.

thanks and regards
sunil kr
askIITian faculty

Sunil Kumar FP
10 years ago
suppose for ex 1M kcl 74.5
volume=1lt
mass of kcl in 1lt=74.5 gm
mass of water in 1 lt of otal volume solution=1000g-74.5gm=.925gm=.925 kg
mass of solvent=.925 kg
molality=1/.925
=1.08
thus in molality wt of the solvent is less than than total volume of the solution in which both are in denominator.because wt of the solute has to be subtracted to get the molality.

thanks and regards
sunil kr
askIITian faculty
varshith reddy
10 years ago
MOLARITY
Indu
10 years ago
Molality. Molarity is moles solute / litres of solution molality is moles solute / kg solvent For molality, you find the mass of solution, then subtract the mass of solute to get the mass of H2O So you use a slightly lower figure for kg of water then you would litres of solution. Thus molality will always be slightly higher. eg 1.00 M NaCl moles NaCl = 1.00 moles Volume solution = 1.00 L molarity = 1.00 M mass NaCl in 1L 1.00 M solution = molar mass x moles = 58.44 g/mol x 1.0 moles= 58.44 g mass H2O (assuming density = 1 g/ml) = 1000 g - 58.44 g = 941.56 g = 0.94156 kg molality = 1.00 mol / 0.94156 kg = 1.06 m The more dilute the solutution, the lower the mass of solute and the more similar kg solvent will be to Litres solution, thus the more dilute the solution the closer molality and molarity will be.